# CBVLM: Training-free Explainable Concept-based Large Vision Language Models for Medical Image Classification

### v1

@@ Abstract
The main challenges limiting the adoption of deep learning-based solutions in medical workflows are the availability of annotated data and the lack of interpretability of such systems. Concept Bottleneck Models (CBMs) tackle the latter by constraining the final disease prediction on a set of predefined and human-interpretable concepts. However, the increased interpretability achieved through these concept-based explanations implies a higher annotation burden. Moreover, if a new concept needs to be added, the whole system needs to be retrained. Inspired by the remarkable performance shown by Large Vision-Language Models (LVLMs) in few-shot settings, we propose a simple, yet effective, methodology, CBVLM, which tackles both of the aforementioned challenges. First, for each concept, we prompt the LVLM to answer if the concept is present in the input image. Then, we ask the LVLM to classify the image based on the previous concept predictions. Moreover, in both stages, we incorporate a retrieval module responsible for selecting the best examples for in-context learning. By grounding the final diagnosis on the predicted concepts, we ensure explainability, and by leveraging the few-shot capabilities of LVLMs, we drastically lower the annotation cost. We validate our approach with extensive experiments across four medical datasets and twelve LVLMs (both generic and medical) and show that CBVLM consistently outperforms CBMs and task-specific supervised methods without requiring any training and using just a few annotated examples. More information on our project page: https://cristianopatricio.github.io/CBVLM/.
@@ 1 Introduction
The field of medical imaging analysis has witnessed significant advancements through the use of supervised deep learning methods. However, their performance greatly depends on the availability and quality of training samples, which are particularly costly to obtain, as clinician expertise is required. The recent rise of Large Vision-Language Models (LVLMs) has also made its way into the medical field, of which Med-LVLMs such as [1, 2, 3, 4] are prominent examples, where pre-trained LVLMs are fine-tuned using medical data. However, despite the improved performance, these Med-LVLMs require even more annotated data. Consequently, a different research direction has emerged, exploring the potential of applying LVLMs to specific downstream tasks with minimal or no supervision, by leveraging their few-shot capabilities [5, 6, 7], particularly the use of in-context learning (ICL) for improving model performance through a short number of examples in the prompt.
In addition to the need for annotated data, another major challenge with deep learning-based medical imaging is the lack of interpretability, which is of the utmost importance in high-stakes decision-making scenarios [8]. One way of introducing interpretability into such frameworks is via concept-based explanations. Concept Bottleneck Models (CBMs) provide this type of explanations by introducing a layer, the “bottleneck”, that predicts the presence or absence of a predefined set of human-interpretable concepts. The final classification prediction is then based on these concepts. Nevertheless, CBMs still require training and annotated data; in fact, the amount of annotated data actually increases since concept annotations are needed on top of the classification labels.
@@
We propose combining the best of both worlds: the interpretability of CBMs and the few-shot abilities of LVLMs to decrease the annotation cost. Our methodology, CBVLM, encompasses a two-stage approach, where, first, the LVLM is prompted to predict the existence/absence of a set of concepts in an image, and, in a second stage, the predicted concepts are incorporated into the prompt such that the LVLM provides a final diagnosis grounded on these predicted concepts. Figure 1 represents this methodology compared to the common approach where an LVLM is simply prompted to perform classification. Moreover, contrary to CBMs, which require retraining the entire model to incorporate a new concept, CBVLM can easily incorporate additional concepts, requiring only two inference steps. We validate CBVLM on four medical datasets encompassing different imaging modalities and using twelve LVLMs, and show that CBVLM achieves comparable performance or even surpasses task-specific supervised methods, doing so without training, with few annotated examples, and without sacrificing interpretability. Although, individually, the techniques used in CBVLM (e.g., ICL) are not novel, this is, to the best of our knowledge, the first work combining them in a novel methodology exploring the capability of LVLMs to provide concept-based explanations for medical image classification.
@@ 2 Related Work
CBMs [14] are based on an encoder-decoder paradigm, where the encoder predicts a set of human-specified concepts from the input image, while the decoder leverages these predicted concepts to generate the final predictions. Due to their inherent interpretability, these models have attracted particular interest in medical image classification. Several CBMs have been proposed for a wide variety of medical tasks (e.g., melanoma diagnosis [15, 16, 17], knee osteoarthritis [14] and chest X-ray classification [18]). Nevertheless, CBMs have well-known limitations: (i)¡ they require concept annotations (image-level or class-level), (ii) they tend to have lower performance compared to traditional (black-box) supervised methods, (iii) they struggle to predict concepts accurately, and (iv) incorporating new concepts implies retraining the whole CBM. These challenges are particularly significant in medical imaging, where data availability is typically limited due to the expertise required for annotations. Consequently, LVLMs represent a promising alternative to traditional CBMs, as they are able to generalize to unseen domains with minimal or no supervision. Moreover, incorporating additional concepts becomes trivial and does not require fine-tuning, as happened with traditional CBMs.
@@ 3 Methodology
Given the task of predicting a target $y\in\mathbb{R}$ from input $x\in\mathbb{R}^{d}$, let $\mathcal{D}=\{(x^{(i)},y^{(i)},c^{(i)})\}_{i=1}^{n}$ be a batch of training samples where $c\in\mathbb{R}^{l}$ is a vector of $l$ concepts. Traditional (black-box) models learn a function $h:\mathbb{R}^{d}\rightarrow\mathbb{R}$ that directly predicts the target $y\in\mathbb{R}$ from the input $x\in\mathbb{R}^{d}$. However, CBMs first map the input $x$ into a set of interpretable concepts $c$ (“bottleneck”) by learning $g:\mathbb{R}^{d}\rightarrow\mathbb{R}^{l}$, and use these concepts to predict the target $y$ using $f:\mathbb{R}^{l}\rightarrow\mathbb{R}$. Therefore, the prediction $\hat{y}=f(g(x))$ is entirely based on the predicted concepts $\hat{c}=g(x)$.
@@ 3.2 Disease Diagnosis
{A brief description of {concept 00}}. (…) {A brief description of {concept $l-1$}. What is the diagnosis that is associated with the following concepts: {predicted concepts} Options: A) class 0 B) class 1 (…) Choose one option. Do not provide additional information. Answer:
@@ 4.1 Datasets
For the Derm7pt [27] dataset, with a total of 827 images, we consider only dermoscopic images of the “nevus” and “melanoma” classes, jointly with the 5 annotated dermoscopic attributes, namely Absent/Typical/Atypical Pigment Network, Absent/Regular/Irregular Streaks, Absent/Regular/Irregular Dots and Globules, Blue-Whitish Veils, and Regression Structures. The SkinCon [28] dataset contains 3,230 images from the Fitzpatrick 17k dataset [31] annotated with 48 clinical concepts. Following previous works, we use the binary labels denoting skin malignancy for the target task and randomly split the dataset into train/val/test with the proportion of 70%, 15%, and 15%, respectively. Additionally, only 22 out of the 48 concepts are used, as the other 26 concepts appear in less than 50 images. From the CORDA [29] dataset, a COVID-19 diagnosis dataset, we use its 1601 chest X-rays. Inspired by the work of Barbano et al. [32], we use a pretrained model on the CheXpert [33] dataset to annotate the samples with a binary label to denote the presence or absence of 14 radiological observations. Of those 14 concepts, 9 are used as they are present in at least 50 images. We also exclude the “No Finding” concept, keeping the remaining 8 concepts. The DDR [30] dataset comprises 13,673 fundus images divided into six Diabetic Retinopathy (DR) levels. In our experiments, we focus on the five-class classification task for DR grading, discarding the images belonging to the ungradable class and considering a subset of 757 images annotated with 4 types of lesions correlated with DR, which we use as concepts. We augment this subset with 743 “No DR” images, resulting in a dataset of 1,500 samples.
@@ 4.4 Retrieval Module
For the retrieval strategy used to perform ICL, image similarity is determined using cosine similarity between features, which are extracted from three distinct encoders, namely CLIP ViT/B-16 [49], MedImageInsight [43] and the (vision) encoder of the corresponding LVLM.
@@ 5 Results and Analysis
This section presents the key results of CBVLM, comparing its performance against CBMs and task-specific supervised methods in terms of BACC and F1-score across two tasks: i) Concept Detection (Figures 3, 4, and 5); and ii) Disease Diagnosis (Figures 6 and 7).
@@ 5.1 Concept Detection Performance
We start by assessing the predictive performance of CBVLM in the task of concept detection. Figure 3 presents the results across different datasets under various $n$-shot settings, $n\in\{0,1,2,4\}$. For comparative analysis, we also report the performance of CBM and CLAT. The results for $n>0$ were obtained considering the best vision encoder, in this case, MedImageInsight, selected after an ablation study across three different vision encoders (see Figure 8).
@@
From the results depicted in Figure 4 (right), in the 00-shot scenario, generic LVLMs consistently outperform medical LVLMs across all datasets. However, in the few-shot setting, medical LVLMs outperform their generic counterparts in all datasets using only one or two demonstration examples ($n\in\{1,2\}$). The stronger performance of generic LVLMs in the zero-shot setting can be attributed to their training on large-scale generic image datasets, allowing for better generalization in this setting. As demonstration examples are introduced into the prompt, medical LVLMs improve, benefiting from their pre-training on specialized medical data, which enhances their ability to generalize in these scenarios.
@@ 5.2 Disease Diagnosis Performance
The results in Figure 7 (right) confirm that generic LVLMs consistently outperform their medical counterparts across all datasets in the scenario without concepts (“0 w/o”). This trend is also observed in the 0-shot scenario with concepts, in general. In the few-shot setting, generic LVLMs exhibit superior performance only in SkinCon in terms of BACC, with a marginal average improvement of 1% across scenarios with $n>0$. Conversely, when averaging across all few-shot scenarios ($n>0$), medical LVLMs consistently outperform generic models across all datasets, achieving F1-score improvements of 25.74%, 14.48%, 4.58%, and 19.81% for Derm7pt, SkinCon, CORDA, and DDR, respectively. This phenomenon can be attributed to two factors: i) the stronger domain knowledge of medical LVLMs, and ii) the initialization of their vision encoders from medical-specific CLIP versions, such as BiomedCLIP [50] for LLaVA-Med, or larger improved CLIP architectures like EVA-CLIP-g [51], used in CheXagent and SkinGPT.
@@
CBVLM surpasses CBMs across all datasets. With the top-performing LVLM, we observe F1-score improvements of 10.88%, 45.71%, 21.04%, and 63.13% for Derm7pt, SkinCon, CORDA, and DDR, respectively, compared to the standard CBM. It is worth recalling that CBVLM achieves these results without any kind of training, whereas CBM needs to be trained for each dataset. When comparing CBVLM with CLAT [41], a method specifically optimized for retinal disease diagnosis, our experiments show that CBVLM outperforms it on the diabetic retinopathy dataset, DDR, in terms of F1-score. Furthermore, CBVLM also outperforms CLAT on the other datasets, achieving an average relative improvement of $\approx$ 6%.
@@
Finally, when compared to black-box approaches, a similar trend is observed. CBVLM outperforms black-box models by 2.15% in BACC for Derm7pt and achieves comparable performance on CORDA, with a slight decrease of -1.07%. In terms of F1-score, the top-performing CBVLM surpasses black-box models by 14.42% on SkinCon and by 1% on DDR. Thus, CBVLM achieves state-of-the-art results on Derm7pt, SkinCon and DDR.
@@ 5.4 General Discussion
Table 2 presents a detailed comparison of our proposed CBVLM with concept-based models, supervised black-box models, and task-specific models. As highlighted in the results, CBVLM consistently achieves either the best or second-best performance across all evaluations. Notably, this is accomplished without any additional training and by leveraging just four examples, enabling concept-based explanations and enhanced interpretability and transparency in the model responses. In summary, CBVLM using CheXagent with 4 demonstration examples emerges as the top-performing combination for Derm7pt. For SkinCon, CBVLM with the generic Open-Flamingo model and 4 demonstration examples achieves the best results in both disease diagnosis and concept detection tasks. For CORDA and DDR, CBVLM with Idefics3 and 4 demonstrations proves to be the best-performing combination.
@@
Dataset Method Disease Diagnosis Concept Detection BACC (%) F1 (%) BACC (%) F1 (%) Derm7pt CBM [14] 74.01 75.66 75.68 67.11 CLAT [41] 79.67 82.98 66.69 54.76 Black-box (ResNet50) 75.33 80.76 - - Black-box (ViT Base) 82.26 82.04 - - Black-box (Task-Specific) [17] 83.20 - - - CBVLM (CheXagent & $4$-shot) 85.35 83.08 70.21 59.55 SkinCon CBM [14] 49.97 46.51 72.81 56.33 CLAT [41] 68.21 86.76 62.64 45.52 Black-box (ResNet50) 73.98 81.49 - - Black-box (ViT Base) 80.25 85.85 - - Black-box (Task-Specific) [46] - 77.80 - - CBVLM (Open-Flamingo & $4$-shot) 85.61 86.21 79.78 57.91 CORDA CBM [14] 57.95 57.85 77.88 66.64 CLAT [41] 76.19 76.71 65.38 54.82 Black-box (ResNet50) 73.17 73.28 - - Black-box (ViT Base) 78.03 77.63 - - Black-box (Task-Specific) [47] 80.00 - - - CBVLM (Idefics3 & $4$-shot) 78.93 78.89 82.87 63.70 DDR CBM [14] 23.62 21.25 59.05 39.06 CLAT [41] 78.87 72.81 76.64 64.53 Black-box (ResNet50) 33.70 57.00 - - Black-box (ViT Base) 39.09 59.37 - - Black-box (Task-Specific) [48] - 83.38 - - CBVLM (Idefics3 & $4$-shot) 66.65 84.38 82.69 75.73

### v2

@@ Abstract
The main challenges limiting the adoption of deep learning-based solutions in medical workflows are the availability of annotated data and the lack of interpretability of such systems. Concept Bottleneck Models (CBMs) tackle the latter by constraining the model output on a set of predefined and human-interpretable concepts. However, the increased interpretability achieved through these concept-based explanations implies a higher annotation burden. Moreover, if a new concept needs to be added, the whole system needs to be retrained. Inspired by the remarkable performance shown by Large Vision-Language Models (LVLMs) in few-shot settings, we propose a simple, yet effective, methodology, CBVLM, which tackles both of the aforementioned challenges. First, for each concept, we prompt the LVLM to answer if the concept is present in the input image. Then, we ask the LVLM to classify the image based on the previous concept predictions. Moreover, in both stages, we incorporate a retrieval module responsible for selecting the best examples for in-context learning. By grounding the final diagnosis on the predicted concepts, we ensure explainability, and by leveraging the few-shot capabilities of LVLMs, we drastically lower the annotation cost. We validate our approach with extensive experiments across four medical datasets and twelve LVLMs (both generic and medical) and show that CBVLM consistently outperforms CBMs and task-specific supervised methods without requiring any training and using just a few annotated examples. More information on our project page: https://cristianopatricio.github.io/CBVLM/.
@@ 1 Introduction
The field of medical imaging analysis has witnessed significant advancements through the use of supervised deep learning methods. However, their performance greatly depends on the availability and quality of training samples, which are particularly costly to obtain, as clinician expertise is required. The recent rise of Large Vision-Language Models (LVLMs) has also made its way into the medical field, of which Med-LVLMs such as [1, 2, 3, 4] are prominent examples, where pre-trained LVLMs are fine-tuned using medical data. However, despite the improved performance, these Med-LVLMs require even more annotated data. Consequently, a different research direction has emerged, exploring the potential of applying LVLMs to specific downstream tasks with minimal or no supervision, by leveraging their few-shot capabilities [5, 6, 7], particularly the use of in-context learning (ICL) for improving model performance through a short number of examples in the prompt. In addition to the need for annotated data, another major challenge with deep learning-based medical imaging is the lack of interpretability, which is of the utmost importance in high-stakes decision-making scenarios [8]. One way of introducing interpretability into such frameworks is via concept-based explanations. Concept Bottleneck Models (CBMs) provide this type of explanations by introducing a layer, the “bottleneck”, that predicts the presence or absence of a predefined set of human-interpretable concepts. The final classification prediction is then based on these concepts. Nevertheless, CBMs still require training and annotated data; in fact, the amount of annotated data actually increases since concept annotations are needed on top of the classification labels. We propose combining the best of both worlds: the interpretability of CBMs and the few-shot abilities of LVLMs to decrease the annotation cost. Our methodology, CBVLM, encompasses a two-stage approach, where, first, the LVLM is prompted to predict the existence/absence of a set of concepts in an image, and, in a second stage, the predicted concepts are incorporated into the prompt such that the LVLM provides a final diagnosis grounded on these predicted concepts. Figure 1 represents this methodology compared to the common approach where an LVLM is simply prompted to perform classification. Moreover, contrary to CBMs, which require retraining the entire model to incorporate a new concept, CBVLM can easily incorporate additional concepts, requiring only two inference steps. We validate CBVLM on four medical datasets encompassing different imaging modalities and using twelve LVLMs, and show that CBVLM achieves comparable performance or even surpasses task-specific supervised methods, doing so without training, with few annotated examples, and without sacrificing interpretability. Although, individually, the techniques used in CBVLM (e.g., ICL) are not novel, this is, to the best of our knowledge, the first work combining them in a novel methodology exploring the capability of LVLMs to provide concept-based explanations for medical image classification.
@@ 2 Related Work
CBMs [14] are based on an encoder-decoder paradigm, where the encoder predicts a set of human-specified concepts from the input image, while the decoder leverages these predicted concepts to generate the final predictions. Due to their inherent interpretability, these models have attracted particular interest in medical image classification. Several CBMs have been proposed for a wide variety of medical tasks (e.g., melanoma diagnosis [15, 16, 17], knee osteoarthritis [14] and chest X-ray classification [18]). Nevertheless, CBMs have well-known limitations: (i) they require concept annotations (image-level or class-level), (ii) they tend to have lower performance compared to traditional (black-box) supervised methods, (iii) they struggle to predict concepts accurately, and (iv) incorporating new concepts implies retraining the whole CBM. These challenges are particularly significant in medical imaging, where data availability is typically limited due to the expertise required for annotations. Consequently, LVLMs represent a promising alternative to traditional CBMs, as they are able to generalize to unseen domains with minimal or no supervision. Moreover, incorporating additional concepts becomes trivial and does not require fine-tuning, as happened with traditional CBMs.
@@ 3 Methodology
Given the task of predicting a target $y\in\mathbb{R}$ from input $x\in\mathbb{R}^{d}$, let $\mathcal{D}=\{(x^{(i)},y^{(i)},c^{(i)})\}_{i=1}^{n}$ be a batch of training samples where $c\in\mathbb{R}^{\{0,1\}^{l}}$ is a vector of $l$ concepts. Traditional (black-box) models learn a function $h:\mathbb{R}^{d}\rightarrow\mathbb{R}$ that directly predicts the target $y\in\mathbb{R}$ from the input $x\in\mathbb{R}^{d}$. However, CBMs first map the input $x$ into a set of interpretable concepts $c$ (“bottleneck”) by learning $g:\mathbb{R}^{d}\rightarrow\mathbb{R}^{l}$, and use these concepts to predict the target $y$ using $f:\mathbb{R}^{l}\rightarrow\mathbb{R}$. Therefore, the prediction $\hat{y}=f(g(x))$ is entirely based on the predicted concepts $\hat{c}=g(x)$.
@@ 3.2 Disease Diagnosis
{A brief description of {concept 0}}. (…) {A brief description of {concept l1l-1}}. What is the diagnosis that is associated with the following concepts: {predicted concepts} Options: A) class 0 B) class 1 (…) Choose one option. Do not provide additional information. Answer:
@@ 4.1 Datasets
For the Derm7pt [27] dataset, with a total of 827 images, we consider only dermoscopic images of the “nevus” and “melanoma” classes, jointly with the 5 annotated dermoscopic attributes, namely Absent/Typical/Atypical Pigment Network, Absent/Regular/Irregular Streaks, Absent/Regular/Irregular Dots and Globules, Blue-Whitish Veils, and Regression Structures. The SkinCon [28] dataset contains 3,230 images from the Fitzpatrick 17k dataset [31] annotated with 48 clinical concepts. Following previous works, we use the binary labels denoting skin malignancy for the target task and randomly split the dataset into train/val/test with the proportion of 70%, 15%, and 15%, respectively. Additionally, only 22 out of the 48 concepts are used, as the other 26 concepts appear in less than 50 images. From the CORDA [29] dataset, a COVID-19 diagnosis dataset, we use its 1601 chest X-rays. Inspired by the work of Barbano et al. [32], we use a pretrained model on the CheXpert [33] dataset to annotate the samples with a binary label to denote the presence or absence of 14 radiological observations. Of those 14 concepts, we exclude the “Fracture” and “Support Devices” concepts, as these are not directly related to the presence or absence of the target class. The “No Finding” concept is implicitly included in the remaining 11 concepts given that if all 11 concepts are 0 it represents a healthy case without any findings. The DDR [30] dataset comprises 13,673 fundus images divided into six Diabetic Retinopathy (DR) levels. In our experiments, we focus on the five-class classification task for DR grading, discarding the images belonging to the ungradable class and considering a subset of 757 images annotated with 4 types of lesions correlated with DR, which we use as concepts. We augment this subset with 743 “No DR” images, resulting in a dataset of 1,500 samples.
@@ 4.4 Retrieval Module
For the retrieval strategy used to perform ICL, image similarity is determined using cosine similarity between features, which are extracted from two distinct encoders, namely MedImageInsight [43] and the (vision) encoder of the corresponding LVLM.
@@ 5 Results and Analysis
This section presents the key results of CBVLM, comparing its performance against CBMs and task-specific supervised methods in terms of BACC and F1-score across two tasks: i) Concept Detection (Figures 3, 4, 5, and 6); and ii) Disease Diagnosis (Figures 7 and 8).
@@ 5.1 Concept Detection Performance
We start by assessing the predictive performance of CBVLM in the task of concept detection. Figure 3 presents the results across different datasets under various $n$-shot settings, $n\in\{0,1,2,4\}$. For comparative analysis, we also report the performance of CBM and CLAT. The results for $n>0$ were obtained considering the best vision encoder, in this case, MedImageInsight, selected after an ablation study with different vision encoders (see Figure 9).
@@
From the results depicted in Figure 4 (right), in the 0-shot scenario, generic LVLMs consistently outperform medical LVLMs across all datasets. However, in the few-shot setting, medical LVLMs outperform their generic counterparts in all datasets using only one or two demonstration examples ($n\in\{1,2\}$). The stronger performance of generic LVLMs in the zero-shot setting can be attributed to their training on large-scale generic image datasets, allowing for better generalization in this setting. As demonstration examples are introduced into the prompt, medical LVLMs improve, benefiting from their pre-training on specialized medical data, which enhances their ability to generalize in these scenarios. However, in the 4-shot setting, generic LVLMs tend to achiever higher concept detection performance than their medical counterparts. This might stem from the fact that generic LVLMs tend to allow for longer prompts and are trained with interleaved image-text data, while medical LVLMs are not.
@@
Figure 5: Per-concept F1-scores across datasets averaged over all (blue), generic (green), and medical (orange) LVLMs, in the 4-shot scenario. Values in parentheses indicate the number of samples in the dataset for the corresponding concept.
@@
Underrepresented concepts are harder to predict
Figure 5 presents a detailed comparison of the per-concept F1-score for each dataset. The reported results, averaged across all, generic, and medical LVLMs in the 4-shot setting, reveal that underrepresented concepts are the most challenging to detect, as expected. For instance, in the CORDA dataset, none of the models correctly predicts the presence of the “pleural other” concept (which only appears in 4 images of the entire dataset). The concepts “atelectasis” and “lung lesion” are also consistently difficult to detect. The full results per-LVLM can be found in the Supplementary Material.
@@ 5.2 Disease Diagnosis Performance
The results in Figure 8 (right) confirm that generic LVLMs consistently outperform their medical counterparts across all datasets in the scenario without concepts (“0 w/o”). This trend is also observed in the 0-shot scenario with concepts, in general. In the few-shot setting, generic LVLMs exhibit superior performance only in SkinCon in terms of BACC, with a marginal average improvement of 1% across scenarios with $n>0$. Conversely, when averaging across all few-shot scenarios ($n>0$), medical LVLMs consistently outperform generic models across all datasets, achieving F1-score improvements of 25.74%, 14.48%, 4.62%, and 19.81% for Derm7pt, SkinCon, CORDA, and DDR, respectively. This phenomenon can be attributed to two factors: i) the stronger domain knowledge of medical LVLMs, and ii) the initialization of their vision encoders from medical-specific CLIP versions, such as BiomedCLIP [49] for LLaVA-Med, or larger improved CLIP architectures like EVA-CLIP-g [50], used in CheXagent and SkinGPT.
@@
CBVLM surpasses CBMs across all datasets. With the top-performing LVLM, we observe F1-score improvements of 7.42%, 39.70%, 7.25%, and 63.13% for Derm7pt, SkinCon, CORDA, and DDR, respectively, compared to the standard CBM. It is worth recalling that CBVLM achieves these results without any kind of training, whereas CBM needs to be trained for each dataset. When comparing CBVLM with CLAT [41], a method specifically optimized for retinal disease diagnosis, our experiments show that CBVLM outperforms it on the diabetic retinopathy dataset, DDR, in terms of F1-score. Furthermore, CBVLM also outperforms CLAT on the other datasets, achieving an average relative improvement of $\approx$ 6%.
@@
Finally, when compared to black-box approaches, a similar trend is observed. CBVLM outperforms task-specific black-box models by 2.15% in BACC for Derm7pt and achieves comparable performance on CORDA, with a slight decrease of -2.32%. In terms of F1 score, the best performing CBVLM exceeds black-box task-specific models by 0.36% on SkinCon and by 1% on DDR. Thus, CBVLM achieves state-of-the-art results on Derm7pt, SkinCon and DDR.
@@ 5.4 General Discussion
Table 2 presents a detailed comparison of our proposed CBVLM with concept-based models, supervised black-box models, and task-specific models. As highlighted in the results, CBVLM consistently achieves either the best or second-best performance across all evaluations. Notably, this is accomplished without any additional training and by leveraging just four examples, enabling concept-based explanations and enhanced interpretability and transparency in the model responses. In summary, CBVLM using CheXagent with 4 demonstration examples emerges as the top-performing combination for Derm7pt. For SkinCon, CBVLM with the generic Open-Flamingo model and 4 demonstration examples achieves the best results in both disease diagnosis and concept detection tasks. For CORDA, the best performance is achieved by CBVLM with CheXagent with 2 demonstration examples. For DDR, CBVLM with Idefics3 and 4 demonstrations proves to be the best-performing combination.
@@
Dataset Method Disease Diagnosis Concept Detection BACC (%) F1 (%) BACC (%) F1 (%) Derm7pt CBM [14] 74.01 75.66 75.68 67.11 CLAT [41] 79.67 82.98 66.69 54.76 Black-box (ResNet50) 75.33 80.76 - - Black-box (ViT Base) 82.26 82.04 - - Black-box (Task-Specific) [17] 83.20 - - - CBVLM (CheXagent & $4$-shot) 85.35 83.08 70.21 59.55 SkinCon CBM [14] 49.97 46.51 72.81 56.33 CLAT [41] 68.21 86.76 62.64 45.52 Black-box (ResNet50) 73.98 81.49 - - Black-box (ViT Base) 80.25 85.85 - - Black-box (Task-Specific) [46] - 77.80 - - CBVLM (Open-Flamingo & $4$-shot) 85.61 86.21 79.78 57.91 CORDA CBM  [14] 70.75 70.98 76.20 61.92 CLAT  [41] 75.64 77.96 70.76 53.06 Black-box (ResNet50) 73.17 73.28 - - Black-box (ViT Base) 78.03 77.63 - - Black-box (Task-Specific) [47] 80.00 - - - CBVLM (CheXagent & $2$-shot) 77.68 78.23 81.38 64.33 DDR CBM [14] 23.62 21.25 59.05 39.06 CLAT [41] 78.87 72.81 76.64 64.53 Black-box (ResNet50) 33.70 57.00 - - Black-box (ViT Base) 39.09 59.37 - - Black-box (Task-Specific) [48] - 83.38 - - CBVLM (Idefics3 & $4$-shot) 66.65 84.38 82.69 75.73